OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4)

OpenStudy (anonymous):

10x-35=33-18x-12 10x+18x=33+35-12 28x=56 x=56/28 x=2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

hey that's fyn...